# Enhancing and Assessing Instruction-Following with Fine-Grained Instruction Variants

### v1

@@ 3.5 Ablation Study
We further conduct ablation study on DeMoRecon-Eval to demonstrate the effectiveness of fine-tuning LLMs on both SFT and DPO losses. Table 3 show the results of fine-tuning LLaMA-2-7B-Chat. In the table, “1603” is the actual number of seed instructions in DeMoRecon datasets. We also exam the trend of performance with increasing seed instruction. As shown in the table, DPO+SFT grants better performance compared with both SFT- and DPO-only fine-tuning, which meet our assumption that learning to follow instructions requires supervision from both generation and preference side.
Furthermore, we observe an increase in performance on DeMoRecon-Eval with the expansion of the seed instruction set, applicable across all fine-tuning methods. While the DPO method achieves noticeable gains more rapidly than the combined DPO+SFT approach, DPO+SFT demonstrates a gradual improvement as the number of seed instructions increases. This trend suggests significant potential for further enhancing nuanced instruction-following capabilities if additional seed instructions are integrated.
@@ 5 Conclusion
In conclusion, we have developed an efficient instruction augmentation method that generates nuanced instruction variants maintaining the same topic and context as the original instructions. Utilizing this method, we constructed DeMoRecon, which comprises five distinct sets tailored for supervised fine-tuning, direct preference optimization, and instruction-following evaluation on fine-grained instruction variants. Our experimental results demonstrate that our combined DPO+SFT tuning strategy, implemented with the DeMoRecon-Aug-DPO and DeMoRecon-Ref-DPO datasets, significantly enhances performance across IFEval, FollowBench, InfoBench, and DeMoRecon-Eval. Our research not only highlights the deficiencies of current popular instruction-tuned LLMs in handling nuanced instruction variants but also provides a robust method and datasets to bridge this gap.
@@
Additionally, the cost of the GPT-4 API poses another limitation. The entire process of augmentation and evaluation depends on GPT-4, a common challenge across many LLM-related studies. Despite this, the costs associated with calling such APIs are decreasing due to the rapid advancement of LLM technologies, making these tools increasingly affordable and accessible.
@@ Appendix A Experimental Settings
Figure 3: An real example from DeMoRecon-Eval. The base model is LLaMA-2-7B-Chat. We show the prediction results of the original model and its DPO tuned version using DeMoRecon-Ref-DPO. Response in red box indicates that GPT-4 judged the response as failing to follow the instruction, while a green box signifies success.
@@ Appendix B Instruction Non-compliance Detection
As DPO seeks to better align LLMs with human preferences, we also explored the relationship between instruction-following accuracy and non-compliance detection. For each instruction in DeMoRecon-Eval, we collect all of its augmented instructions and then paired each augmented instruction with every GPT-4 response from those augmented instructions. Each instruction-response pair was labeled ‘Yes’ if the response ground truth of the instruction, and ‘No’ otherwise. This methodology ensured a minimum of two sub-instruction differences between the correct instruction and any incorrect responses, enhancing the detection challenge.
@@
Table 5: Experimental results on DeMoRecon-Det .
Model DeMoRecon-Eval Ori. SFT DPO+SFT Plain Aug. Ref. Aug. Ref. Llama2-7B-Chat 53.43 53.31 52.23 53.55 55.59 58.04 Qwen1.5-7B-Chat 71.18 58.31 73.63 58.75 78.55 73.43 Mistral-7B-Instruct 65.66 71.89 68.97 69.51 80.17 83.28 Llama2-13B-Chat 65.02 69.85 68.55 69.26 77.23 79.66 Qwen1.5-14B-Chat 70.86 69.95 72.60 73.5 86.37 86.76 Orion-14B-Chat 63.82 62.08 65.66 63.60 77.16 79.19
Table 6: Results on DeMoRecon-Det using difference fine-tuning method and number of original instructions. The bash model is LLaMA-2-7B-Chat
@@ Appendix C Case Study
Figure 3 presents a real-case example from the DeMoRecon-Eval set, showcasing an instruction and its three variants, each illustrating typical errors made by LLMs in following instructions.
@@ Appendix D Full Experimental Results
Model FollowBench (HSR) InfoBench Level 1 Level 2 Level 3 Level 4 Level 5 Avg. Easy Hard Average LLaMA-2-7B-Chat Original 57.68 51.96 38.16 36.02 31.67 43.10 79.57 76.60 77.51 SFT Plain 61.55 51.51 41.85 38.67 28.86 44.49 84.20 76.28 78.71 Augmentation 57.31 50.91 44.07 36.02 31.78 44.02 81.88 77.31 78.71 Reference 59.74 51.68 42.54 36.58 25.89 43.28 81.88 73.59 76.13 SFT+DPO Augmentation 61.23 55.18 43.31 36.93 31.17 45.56 82.61 78.91 80.04 Reference 63.29 56.38 44.41 34.95 33.00 46.41 82.17 80.58 81.07 Qwen1.5-7B-Chat Original 74.35 58.80 52.85 44.66 34.33 53.00 84.49 78.85 80.58 SFT Plain 70.42 61.57 51.77 42.66 40.37 53.36 83.62 77.37 79.29 Augmentation 69.17 58.55 48.37 43.33 36.03 51.09 84.93 77.47 79.73 Reference 73.75 57.34 55.23 42.58 34.96 52.77 85.07 75.90 78.71 SFT+DPO Augmentation 70.22 59.69 50.72 37.89 38.29 51.36 84.93 79.42 81.11 Reference 69.64 61.57 57.36 47.98 37.75 54.86 86.09 78.78 81.02 Mistral-7B-Instruct Original 61.70 52.29 40.92 34.18 23.67 42.55 85.22 76.47 79.16 SFT Plain 61.52 50.44 47.09 39.82 33.93 46.56 84.64 75.26 78.13 Augmentation 59.63 55.77 47.18 40.40 32.94 47.18 84.35 75.71 78.36 Reference 61.06 53.95 48.14 37.93 30.77 46.37 83.77 76.35 78.62 SFT+DPO Augmentation 60.46 56.50 49.38 45.76 36.02 49.63 85.65 76.92 79.60 Reference 59.27 56.93 48.45 41.93 31.80 47.68 84.64 78.78 80.58 LLaMA-2-13B-Chat Original 56.36 54.74 45.31 37.80 30.49 44.94 83.91 79.81 81.07 SFT Plain 60.28 53.03 42.59 41.94 34.96 46.56 83.91 79.68 80.98 Augmentation 59.95 51.67 43.28 39.11 35.26 45.85 84.49 79.04 80.71 Reference 59.72 54.08 45.35 39.30 35.52 46.79 84.93 79.49 81.16 SFT+DPO Augmentation 64.79 53.91 46.35 41.39 29.73 47.23 86.38 81.28 82.84 Reference 61.22 55.40 47.88 40.51 35.33 48.07 86.09 80.17 82.36 Qwen1.5-14B-Chat Original 76.99 60.19 57.81 45.31 36.32 55.32 84.49 81.54 82.44 SFT Plain 67.87 55.85 49.77 43.01 33.95 50.09 83.48 79.55 80.76 Augmentation 67.17 58.04 52.33 47.05 35.69 52.05 82.75 79.55 80.53 Reference 71.99 55.98 54.59 43.17 33.92 51.93 84.20 77.05 79.24 SFT+DPO Augmentation 79.81 65.28 59.00 48.94 37.06 58.02 86.09 82.12 83.33 Reference 77.42 60.37 51.99 49.45 42.00 56.25 85.36 82.31 83.24 Orion-14B-Chat Original 69.42 57.81 41.97 27.65 23.73 44.12 83.33 73.33 76.40 SFT Plain 66.54 55.87 45.43 45.13 31.09 48.81 82.03 73.46 76.09 Augmentation 69.08 55.59 48.51 42.53 30.13 49.17 83.19 74.36 77.07 Reference 67.32 56.72 45.75 40.93 32.83 48.71 82.32 71.22 74.62 SFT+DPO Augmentation 73.78 60.85 52.99 45.18 35.29 53.62 84.35 76.28 78.76 Reference 73.44 61.25 53.48 43.54 38.27 54.00 85.07 75.90 78.71
@@ E.4 Collect Response with Inference
E.5 DeMoRecon-Eval Evaluation
@@
E.6 DeMoRecon-Det Evaluation
We employ a specific prompt template to construct DeMoRecon-Det, which is designed to evaluate the LLM’s ability to detect instruction non-compliance. Each pair of instruction-response data is inserted into this template, serving as the prompt for LLMs to assess whether the response adheres to the given instruction.

### v3

@@ 3.5 Ablation Study
We further conduct ablation study on FGIV-Eval to demonstrate the effectiveness of fine-tuning LLMs on both SFT and DPO losses. Table 3 show the results of fine-tuning LLaMA-2-7B-Chat. In the table, “1603” is the actual number of seed instructions in FGIV datasets. We also exam the trend of performance with increasing seed instruction. As shown in the table, DPO+SFT grants better performance compared with both SFT- and DPO-only fine-tuning, which meet our assumption that learning to follow instructions requires supervision from both generation and preference side.
Furthermore, we observe an increase in performance on FGIV-Eval with the expansion of the seed instruction set, applicable across all fine-tuning methods. While the DPO method achieves noticeable gains more rapidly than the combined DPO+SFT approach, DPO+SFT demonstrates a gradual improvement as the number of seed instructions increases. This trend suggests significant potential for further enhancing nuanced instruction-following capabilities if additional seed instructions are integrated.
@@ 5 Conclusion
In conclusion, we have developed an efficient instruction augmentation method that generates nuanced instruction variants maintaining the same topic and context as the original instructions. Utilizing this method, we constructed FGIV, which comprises three distinct sets tailored for supervised fine-tuning, direct preference optimization, and instruction-following evaluation on fine-grained instruction variants. Our experimental results demonstrate that our combined DPO+SFT tuning strategy, implemented with the FGIV-A and FGIV-R datasets, significantly enhances performance across IFEval, FollowBench, InfoBench, and FGIV-Eval. Our research not only highlights the deficiencies of current popular instruction-tuned LLMs in handling nuanced instruction variants but also provides a robust method and datasets to bridge this gap.
@@
Additionally, the cost of the GPT-4 API presents another limitation. The entire augmentation process totaled approximately 500 USD. The process of augmentation and evaluation heavily relies on GPT-4, a common challenge in many LLM-related studies. However, the costs associated with using such APIs are decreasing due to the rapid advancements in LLM technologies, making these tools increasingly affordable and accessible.
@@ A.2 Experimental Settings
Figure 3: An real example from FGIV-Eval. The base model is LLaMA-2-7B-Chat. We show the prediction results of the original model and its DPO tuned version using FGIV-R. Response in red box indicates that GPT-4 judged the response as failing to follow the instruction, while a green box signifies success.
@@ Appendix B Instruction Non-compliance Detection
As DPO seeks to better align LLMs with human preferences, we also explored the relationship between instruction-following accuracy and non-compliance detection. For each instruction in FGIV-Eval, we collect all of its augmented instructions and then paired each augmented instruction with every GPT-4 response from those augmented instructions. Each instruction-response pair was labeled ‘Yes’ if the response ground truth of the instruction, and ‘No’ otherwise. This methodology ensured a minimum of two sub-instruction differences between the correct instruction and any incorrect responses, enhancing the detection challenge.
@@
Model FGIV-Eval Ori. SFT DPO+SFT Plain Aug. Ref. Aug. Ref. Llama2-7B-Chat 53.43 53.31 52.23 53.55 55.59 58.04 Qwen1.5-7B-Chat 71.18 58.31 73.63 58.75 78.55 73.43 Mistral-7B-Instruct 65.66 71.89 68.97 69.51 80.17 83.28 Llama2-13B-Chat 65.02 69.85 68.55 69.26 77.23 79.66 Qwen1.5-14B-Chat 70.86 69.95 72.60 73.5 86.37 86.76 Orion-14B-Chat 63.82 62.08 65.66 63.60 77.16 79.19
Table 5: Experimental results on FGIV-Det.
@@
Table 6: Results on FGIV-Det using difference fine-tuning method and number of original instructions. The bash model is LLaMA-2-7B-Chat
@@ Appendix C Case Study
Figure 3 presents a real-case example from the FGIV-Eval set, showcasing an instruction and its three variants, each illustrating typical errors made by LLMs in following instructions.
@@ Appendix D Full Experimental Results
Model FollowBench (HSR) InfoBench Level 1 Level 2 Level 3 Level 4 Level 5 Avg. Easy Hard Average LLaMA-2-7B-Chat Chat 57.68 51.96 38.16 36.02 31.67 43.1 79.57 76.6 77.51 Seed 61.55 51.51 41.85 38.67 28.86 44.49 84.2 76.28 78.71 FGIV-A 61.23 55.18 43.31 36.93 31.17 45.56 82.61 78.91 80.04 FGIV-R 63.29 56.38 44.41 34.95 33 46.41 82.17 80.58 81.07 Qwen1.5-7B-Chat Chat 74.35 58.8 52.85 44.66 34.33 53 84.49 78.85 80.58 Plain 70.42 61.57 51.77 42.66 40.37 53.36 83.62 77.37 79.29 FGIV-A 70.22 59.69 50.72 37.89 38.29 51.36 84.93 79.42 81.11 FGIV-R 69.64 61.57 57.36 47.98 37.75 54.86 86.09 78.78 81.02 Mistral-7B-Instruct Chat 61.7 52.29 40.92 34.18 23.67 42.55 85.22 76.47 79.16 Plain 61.52 50.44 47.09 39.82 33.93 46.56 84.64 75.26 78.13 FGIV-A 60.46 56.5 49.38 45.76 36.02 49.63 85.65 76.92 79.6 FGIV-R 59.27 56.93 48.45 41.93 31.8 47.68 84.64 78.78 80.58 LLaMA-2-13B-Chat Chat 56.36 54.74 45.31 37.8 30.49 44.94 83.91 79.81 81.07 Plain 60.28 53.03 42.59 41.94 34.96 46.56 83.91 79.68 80.98 FGIV-A 64.79 53.91 46.35 41.39 29.73 47.23 86.38 81.28 82.84 FGIV-R 61.22 55.4 47.88 40.51 35.33 48.07 86.09 80.17 82.36 Qwen1.5-14B-Chat Chat 76.99 60.19 57.81 45.31 36.32 55.32 84.49 81.54 82.44 Plain 67.87 55.85 49.77 43.01 33.95 50.09 83.48 79.55 80.76 FGIV-A 79.81 65.28 59 48.94 37.06 58.02 86.09 82.12 83.33 FGIV-R 77.42 60.37 51.99 49.45 42 56.25 85.36 82.31 83.24 Orion-14B-Chat Chat 69.42 57.81 41.97 27.65 23.73 44.12 83.33 73.33 76.4 Plain 66.54 55.87 45.43 45.13 31.09 48.81 82.03 73.46 76.09 FGIV-A 73.78 60.85 52.99 45.18 35.29 53.62 84.35 76.28 78.76 FGIV-R 73.44 61.25 53.48 43.54 38.27 54 85.07 75.9 78.71
@@ E.4 Collect Response with Inference
E.5 FGIV-Eval Evaluation
@@
E.6 FGIV-Det Evaluation
We employ a specific prompt template to construct FGIV-Det, which is designed to evaluate the LLM’s ability to detect instruction non-compliance. Each pair of instruction-response data is inserted into this template, serving as the prompt for LLMs to assess whether the response adheres to the given instruction.